OpenStudy (anonymous):

Find the surface area and volume of a sphere with a radius of 7cm. Use pi equals 3.14 and round to the nearest hundredth

OpenStudy (anonymous):

surface area = 4*pi*r^2 volume = (4/3) * pi * r^3

OpenStudy (anonymous):

use calculator to find 4*3.14*(7)^2; that becomes 616 sq. cm

OpenStudy (anonymous):

plug the values in and use your calculator to get results

OpenStudy (anonymous):

for volume (4/3)*3.14*(7)^3 cc